# Mini Diffuser: Train a Multi-Task Diffusion Policy on RLBench-18 in One Day with One GPU

### v1

@@ Abstract
We present a method for training multi-task vision-language robotic diffusion policies that reduces training time and memory usage by an order of magnitude. This improvement arises from a previously underexplored distinction between action diffusion and the image diffusion techniques that inspired it: image generation targets are high-dimensional, while robot actions lie in a much lower-dimensional space. Meanwhile, the vision-language conditions for action generation remain high-dimensional. Our approach, Mini-Diffuser, exploits this asymmetry by introducing Level-2 minibatching, which pairs multiple noised action samples with each vision-language condition, instead of the conventional one-to-one sampling strategy. To support this batching scheme, we introduce architectural adaptations to the diffusion transformer that prevent information leakage across samples while maintaining full conditioning access. In RLBench simulations, Mini-Diffuser achieves 95% of the performance of state-of-the-art multi-task diffusion policies, while using only 5% of the training time and 7% of the memory. Real-world experiments further validate that Mini-Diffuser preserves the key strengths of diffusion-based policies, including the ability to model multimodal action distributions and produce behavior conditioned on diverse perceptual inputs. Code available at github.com/utomm/mini-diffuse-actor, along with training logs.
@@ 4.4 Design Choices
We propose two variants of our Mini-Diffuser architecture that integrate the three non-invasive building blocks—masked attention, local kernel-based query, and FiLM modulation—into transformer-style diffusion models.
@@ 5.2 Ablation Study
Figure 4: Policy Learning Efficiency. The y-axis shows the proportion of generated actions with errors below specified thresholds (e.g., $<3$cm or $<$3°), indicating successful diffusion. Increasing the Level-2 batch size accelerates convergence at the same number of gradient steps (x-axis).
@@
Table 3 also evaluates other design decisions. When we revert Mini-Diffuser to match the original 3D Diffuse Actor—by replacing the PTv3 backbone and action head—performance remains comparable, but training time increases by 40% and memory usage by 30%. Removing the 3D RoPE module leads to severe overfitting, decreasing success by 17.5%, indicating the critical role of relative spatial encoding. Local kernel-based queries contribute a modest 1% improvement in success rate, but we retain them as they help stabilize early training and add negligible computational cost.
@@ 6 Discussion and Conclusion
References are important to the reader; therefore, each citation must be complete and correct. If at all possible, references should be commonly available publications.

### v2

# Mini Diffuser: Fast Multi-task Diffusion Policy Training Using Two-level Mini-batches

@@ Abstract
We present a method that reduces, by an order of magnitude, the time and memory needed to train multi-task vision-language robotic diffusion policies. This improvement arises from a previously underexplored distinction between action diffusion and the image diffusion techniques that inspired it: In image generation, the target is high-dimensional. By contrast, in action generation, the dimensionality of the target is comparatively small, and only the image condition is high-dimensional. Our approach, Mini Diffuser, exploits this asymmetry by introducing two-level minibatching, which pairs multiple noised action samples with each vision-language condition, instead of the conventional one-to-one sampling strategy. To support this batching scheme, we introduce architectural adaptations to the diffusion transformer that prevent information leakage across samples while maintaining full conditioning access. In RLBench simulations, Mini-Diffuser achieves 95% of the performance of state-of-the-art multi-task diffusion policies, while using only 5% of the training time and 7% of the memory. Real-world experiments further validate that Mini-Diffuser preserves the key strengths of diffusion-based policies, including the ability to model multimodal action distributions and produce behavior conditioned on diverse perceptual inputs. Code available at: mini-diffuse-actor.github.io along with videos and training logs.
@@ 4.4 Design Choices
We explore two variants that integrate our three non-invasive building blocks—masked attention, local kernel–based query, and FiLM modulation—into transformer-style diffusion models. The first variant serves as a lightweight modified baseline which we used for ablation, while the second, which we refer to as the Mini-Diffuser, represents our fully optimized architecture.
@@ 5.2 Ablation Study
Figure 4: Policy Learning Efficiency. The y-axis shows the proportion of generated actions with errors below specified thresholds (e.g., $<1$cm or $<$3°), indicating successful diffusion. Increasing the Level-2 batch size accelerates convergence at the same number of gradient steps (x-axis).
@@
Table 3 also evaluates other design decisions. When we revert Mini-Diffuser to match the original 3D Diffuse Actor—by replacing the PTv3 backbone and action head—performance remains comparable, but training time increases by 18% and memory usage by 44%. Removing the 3D RoPE module leads to severe overfitting, decreasing success by 13.4%, indicating the critical role of relative spatial encoding. Local kernel-based queries contribute a modest 1% improvement in success rate, but we retain them as they help stabilize early training and add negligible computational cost.
@@ 6 Discussion and Conclusion
Acknowledgments and Disclosure of Funding
This work is supported by Interne Fondsen KU Leuven/Internal Funds KU Leuven (C2E/24/034). The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government. Part of calculations were also run on the Euler cluster of ETH Zürich